# Learning Multi-Agent Loco-Manipulation for Long-Horizon Quadrupedal Pushing

### v2

@@ I Introduction
  • We propose a hierarchical MARL framework with three hierarchies that can handle long-horizon collaborative quadrupedal pushing in an environments with obstacles.
@@ II-B Multi-Agent Collaborative Manipulation
Optimization-based methods have proven effective in multi-agent collaborative manipulation across various robotic embodiments, such as mobile robots [36, 37, 38, 39], robotic arms [40], quadrotors [41], and six-legged robots [42]. Some works explore utilizing Model Predictive Control (MPC) to achieve cooperative locomotion for multiple quadrupedal robots holonomically constrained to one another [43, 44, 45] or collaborative loco-manipulation with objects rigidly attached to each robotic hand [46]. However, these approaches might lack generalizability to more typical scenarios due to their reliance on specific inter-robot connections. The work in [22] introduces a hierarchical adaptive control method enabling multiple quadrupeds to cooperatively push an object with unknown properties along a predetermined path, though the robots are constrained to use their head to push the objects.
@@ II-C Hierarchical Reinforcement Learning
Hierarchical reinforcement learning (HRL) is often utilized to tackle challenging long-horizon decision making problems. In HRL methods, the high-level policy usually learns to set subgoals for the low-level [56, 57, 58, 59], or learns to combine and chain behavior primitives [60, 50, 61]. In multi-agent settings, the high-level policy in hierarchical MARL generates goals or commands in a decentralized manner [24, 25, 62], or through a centralized controller [63, 64, 23, 65]. Meanwhile, many learning-based controllers for legged robots follow a hierarchical structure, where a high-level RL policy provides intermediate commands to the low-level controllers, such as torso velocities [21, 24], foot landing positions [66, 67], target poses [68, 69], gait timing [70, 71] or a combined of several [72, 73]. In our approach, we use a centralized high-level controller to propose a shared object-centric goal for all robots, while decentralized mid-level controllers send torso velocity commands to each robot’s low-level policy.
@@ III-D High-Level Controller
The high-level controller is composed of two elements, a RRT planner $\mathcal{P}:\mathcal{M}\times\mathcal{G}\rightarrow\mathcal{T}$ and a centralized adaptive policy $\pi^{h}_{\theta}:\mathcal{M}\times\mathcal{T}\times\mathcal{S}_{\text{object}}\times\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\mathcal{S}_{N}\rightarrow\mathcal{A}^{h}$, where $\mathcal{M}$ represents the map information space, $\mathcal{G}$ represents the goal space of the target object, $T$ represents the trajectory space of the RRT planner, $\mathcal{S}_{\text{object}}$ denotes the object state space, $\mathcal{S}^{i}$ is the state space of robot $i$, and $A^{h}$ is the action space of the high-level adaptive policy.
@@ IV-A2 Baselines
Mid-Level+Low-Level with Fine-Tuned Rewards (M+L FT) follows the same policy architecture as M+L with our fine-tuned reward functions. We maintain a complete set of rewards, $r^{m}_{\text{task}}$, $r^{m}_{\text{penalty}}$, and $r^{m}_{\text{heuristic}}$ mentioned in Sec. III-E1. This baseline further enhances the methods in [25] and [24], while also offering an ablation study of our approach without the high-level controller..
@@ IV-B3 Scalability Analysis
To evaluate the scalability of the proposed method, we experiment with different numbers of robots to push the cylinder described in Sec. IV-A1. As shown in Fig. 5, both the success rate and the completion time improve significantly as the number of robots increases. However, once the number of robots reaches four, the success rate decreases compared to that with three robots. This decline could be due to an increased risk of collisions, which causes robots to maintain greater distances from each other, which leads to less coordinated behaviors.
@@ IV-D1 Pushing T
As shown in Figure 6(a), the robots effectively control the turning of the T-shaped block, and successfully push the T-block to the target position. Due to the smaller size of the T-shaped block, finding appropriate pushing points is crucial for the task. We observe that the two robots consistently push from the two ends of the block to apply a larger contact surface, which ensures the application of continuous forward force while also maintaining directional control. Additionally, we find that the $x$ direction velocity commands output by the mid-level policy almost always reach the upper bound of 0.5 m///s. This outcome is expected because the robot is trained to finish the task within shorter time. Once the robot identifies the correct pushing points, its motion becomes predominantly forward, with turning primarily achieved through adjustments in its yaw. Since the target is located in the positive $y$ direction, we observe that the yaw command is initially positive and gradually returns to zero or negative after the turn is completed to straighten the robots. This process is complemented by minor adjustment on $y$ direction velocity commands, which help maintain good contact positions for both robots when interacting with the smaller T-shaped object.
@@ V-A Training Details
of thein simulation approach describ. For detailed implementation information, we direct readers to this prior workThe training and setups for low-level locomotion policy follow the scheme introduced in [73], and we refer our readers to the work for details of implementation.

### v3

@@ I Introduction
  • We propose a hierarchical MARL framework with three hierarchies that can handle long-horizon collaborative quadrupedal pushing in an environment with obstacles.
@@ II-B Multi-Agent Collaborative Manipulation
Optimization-based methods have proven effective in multi-agent collaborative manipulation across various robotic embodiments, such as mobile robots [36, 37, 38, 39], robotic arms [40], quadrotors [41], and six-legged robots [42]. Some works explore utilizing Model Predictive Control (MPC) to achieve cooperative locomotion for multiple quadrupedal robots holonomically constrained to one another [43, 44, 45] or collaborative loco-manipulation with objects rigidly attached to each robotic hand [46]. However, these approaches might lack generalizability to more typical scenarios due to their reliance on specific inter-robot connections. The work in [22] introduces a hierarchical adaptive control method enabling multiple quadrupeds to cooperatively push an object with unknown properties along a predetermined path, though the robots are constrained to use their head to push the objects, while our task requires more diverse contact patterns.
@@ II-C Hierarchical Reinforcement Learning
Hierarchical reinforcement learning (HRL) is often utilized to tackle challenging long-horizon decision making problems. In HRL methods, the high-level policy usually learns to set subgoals for the low-level [56, 57, 58, 59], or learns to combine and chain behavior primitives [60, 50, 61]. In multi-agent settings, the high-level policy in hierarchical MARL generates goals or commands in a decentralized manner [24, 25, 62], or through a centralized controller [63, 64, 23, 65]. Meanwhile, many learning-based controllers for legged robots follow a hierarchical structure, where a high-level RL policy provides intermediate commands to the low-level controllers, such as torso velocities [21, 24], foot landing positions [66, 67], target poses [68, 69], gait timing [70, 71] or a combination of several [72, 73]. In our approach, we use a centralized high-level controller to propose a shared object-centric goal for all robots, while decentralized mid-level controllers send torso velocity commands to each robot’s low-level policy.
@@ III-D High-Level Controller
The high-level controller is composed of two elements, a RRT planner $\mathcal{P}:\mathcal{M}\times\mathcal{G}\rightarrow\mathcal{T}$ and a centralized adaptive policy $\pi^{h}_{\theta}:\mathcal{M}\times\mathcal{T}\times\mathcal{S}_{\text{object}}\times\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\mathcal{S}_{N}\rightarrow\mathcal{A}^{h}$, where $\mathcal{M}$ represents the map information space, $\mathcal{G}$ represents the goal space of the target object, $\mathcal{T}$ represents the trajectory space of the RRT planner, $\mathcal{S}_{\text{object}}$ denotes the object state space, $\mathcal{S}_{i}$ is the state space of robot $i$, and $A^{h}$ is the action space of the high-level adaptive policy.
@@ IV-A2 Baselines
Mid-Level+Low-Level with Fine-Tuned Rewards (M+L FT) follows the same policy architecture as M+L with our fine-tuned reward functions. We maintain a complete set of rewards, $r^{m}_{\text{task}}$, $r^{m}_{\text{penalty}}$, and $r^{m}_{\text{heuristic}}$ mentioned in Sec. III-E1. This baseline further enhances the methods in [25] and [24], while also offering an ablation study of our approach without the high-level controller.
@@ IV-B3 Scalability Analysis
To evaluate the scalability of the proposed method, we experiment with different numbers of robots to push the cylinder described in Sec. IV-A1. As shown in Fig. 5, a single robot struggles with cylinder pushing, while two robots significantly improve success rate with reasonable completion time. Three robots further enhance both metrics, achieving the best performance. However, with four robots, performance slightly declines due to the cylinder’s limited effective contact surface, increasing collision risk and pushing interference (see attached video).
@@ IV-D1 Pushing T
As shown in Figure 6(a), the robots effectively control the turning of the T-shaped block, and successfully push the T-block to the target position. Due to the smaller size of the T-shaped block, finding appropriate pushing points is crucial for the task. We observe that the two robots consistently push from the two ends of the block to apply a larger contact surface, which ensures the application of continuous forward force while also maintaining directional control. Additionally, we find that the $x$ direction velocity commands output by the mid-level policy almost always reach the upper bound of 0.5 m///s. This outcome is expected because the robot is trained to finish the task within shorter time. Once the robot identifies the correct pushing points, its motion becomes predominantly forward, with turning primarily achieved through adjustments in its yaw. Since the target is located in the positive $y$ direction, we observe that the yaw command is initially positive and gradually returns to zero or negative after the turn is completed to straighten the robots. This process is complemented by minor adjustment on the $y$ direction velocity commands, which helps maintain good contact positions for both robots when interacting with the smaller T-shaped object.
@@ V-A Training Details
The training and setup of the low-level locomotion policy in simulation follow the approach described in [73]. For detailed implementation information, we direct readers to this prior work.